# Weighing the Axion with Muon Haloscopy

### 08 Dec 2023 v36

@@ Abstract
Recent measurements of muon spin precession confirm a long-standing tension with the Standard Model of particle physics. We argue that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument yields a percent level prediction for the mass of the axion provided that dark matter is made of axions. An analysis of charge asymmetry in kaon decays suggests that at least in the local halo dark matter is made of axions and that axions from the local halo are responsible for the observed violation by these reactions of the combined charge conjugation and spatial inversion symmetry operation. A tabletop experiment to directly detect dark matter in the form of axions with the predicted mass is proposed.
@@
$A_{M}$ $=$ $M_{L}^{-}\overline{M}_{R}^{-}-M_{R}^{-}\overline{M}_{L}^{-}-M_{L}^{+}\overline{M}_{R}^{+}+M_{R}^{+}\overline{M}_{L}^{+}$ (1)
@@
Next, we suppose that axions $A_{M}$ of each kind and photons $\gamma_{H}$ of each helicity formed in equal numbers in the early universe prior to baryogenesis. This fixes the ratio $n_{A}/n_{\gamma}=|M|/|H|=6$ of axions to photons and gives the axion mass $m_{A}$ ($\hbar=c=1$) [8]:
$m_{A}=2.70~{}kT_{\gamma}\left(\frac{n_{A}}{n_{\gamma}}\right)^{-1}\frac{\Omega_{A}h^{2}}{\Omega_{\gamma}h^{2}}=(0.508\pm 0.004)~{}{\rm eV},$ (2)
where, $T_{\gamma}=(2.7255\pm 0.0006)~{}{\rm K}$ is the present photon temperature [9], $\Omega_{\gamma}h^{2}=2.473\times 10^{-5}$ is the photon energy density parameter and $\Omega_{A}h^{2}=0.11882\pm 0.00086$ is that of dark matter [10].
Local Dark Matter Halo—The energy density of axions in the local dark matter halo of the galaxy $\rho_{A}$ can be estimated using charge asymmetry of kaon decays [11]:
$A_{L}(e)$ $=$ $-2\sqrt{2}\xi_{A\pi}p_{K}\tilde{\phi}_{A}(q_{A})$ (3)
$=$ $4\pi C_{A\pi}\frac{f_{\pi}\sqrt{\rho_{A}hc}}{m_{A}f_{A}}L_{K},$
where, the axion amplitude $\tilde{\phi}_{A}(q_{A})$ at momentum $q_{A}$ changes sign under the combination of charge conjugation and space inversion probed by $A_{L}(e)$ to create the observed charge asymmetry [6]. The mixing angle $\xi_{A\pi}=-C_{A\pi}f_{\pi}/f_{A}$ between the axion $A$ and the neutral pion $\pi^{0}$ can be expressed in terms of the pion decay constant $f_{\pi}=(92.32\pm 0.09)~{}{\rm MeV}$ [12], the axion decay constant $f_{A}$ [13], and the light quark masses using $C_{A\pi}=(m_{d}-m_{u})/(2m_{u}+2m_{d})=(0.173\pm 0.008)$ [14]. Here, the kaon beam has momentum $p_{K}$ and decays in-flight within a detector of length $L_{K}$.
@@
In contrast to the local axion energy density, the axion mass $m_{A}$ and decay constant $f_{A}$ only enter the kaon charge asymmetry through the product $m_{A}f_{A}=\sqrt{\chi_{\rm QCD}}=(5.69\pm 0.05)\times 10^{-3}~{}{\rm GeV^{2}}$ which is fixed by the topological susceptibility $\chi_{\rm QCD}$ of the quantum chromodynamic vacuum [15]. Using the kaon parameters $L_{K}=500~{}{\rm cm}$ and $A_{L}(e)=(3.32\pm 0.08)\times 10^{-3}$ from the current best charge asymmetry measurement [16, 17], the local axion energy density $\rho_{A}=(0.28\pm 0.03)~{}{\rm GeV/cm^{3}}$ that we find from Eq. (3) agrees with the best available estimate $\rho_{\rm DM}=(0.30\pm 0.03)~{}{\rm GeV/cm^{3}}$ of the energy density of the local dark matter halo [18]. This agreement suggests that axions saturate the local dark matter halo and that dark matter from the local halo is the source of the charge asymmetry observed in kaon decays.
Muon Haloscopy—Figure 1 shows how axions in the local dark matter halo shift muon spin precession [3, 4, 5]. Acting as a background field $\tilde{\phi}_{A}(q_{A})$ with on-shell four-momentum $q_{A}^{2}=m_{A}^{2}\ll m_{\mu}^{2}$, the axion $A$ couples to the virtual photon dressing the vertex of the positive muon $\mu^{+}$ and the magnetic field $B$. The axion-photon coupling of strength $g_{A\gamma\gamma}=(0.68\pm 0.02)\times 10^{-10}~{}{\rm GeV^{-1}}$ [19] changes the photon’s Green’s function [20, 21]
$\Delta D_{+}(x,x^{\prime})=m_{A}^{2}g_{A\gamma\gamma}\tilde{\phi}(q_{A})D_{+}(x,x^{\prime}),$ (4)
@@
$\frac{\Delta a_{\mu}}{a_{\mu}}$ $=$ $m_{A}^{2}g_{A\gamma\gamma}\tilde{\phi}_{A}(q_{A})$ (5)
$=$ ±$\pm m_{A}g_{A\gamma\gamma}\sqrt{\rho_{A}N_{\mu}V_{\mu}T_{\mu}}$
where, the axion field strength $|\tilde{\phi}_{A}(q_{A})|^{2}\approx\phi_{A}^{2}N_{\mu}V_{\mu}T_{\mu}$ is fixed by the axion energy density $\rho_{A}=\phi_{A}^{2}m_{A}^{2}$, the volume $V_{\mu}$ of the muon beam, the lifetime $T_{\mu}$ of the muon in the axion rest-frame, and the average number of muons $N_{\mu}$ during spin precession.
@@
$\frac{\Delta a_{\mu}}{a_{\mu}}$ $=$ $(165\pm 10)\times 10^{-11}\left(\frac{N_{\mu}}{1930}\right)^{1/2}\left(\frac{\sigma_{x}}{1.78~{}{\rm cm}}\right)^{1/2}$
@@
where, $\sigma_{x}$ is the width of the muon beam, $\sigma_{y}$ is its height, and $R$ is the radius of the storage ring. This has roughly the right size to resolve the $3.4~{}\sigma$ tension between the spin precession experiments (Exp) [2, 3, 4] and standard model calculations (SM) [25, 26, 27, 28, 29, 30]
@@
where, we have used the recently emerged consensus value $a_{\mu}^{\rm HLO}({\rm SM})=(704.1\pm 3.0)\times 10^{-10}$ [31] for the leading order hadronic contribution to the muon spin precession frequency within the Standard Model [32]. Tabletop Dark Matter Detector— Axions from the local dark matter halo of the galaxy can be seen in the mid-infrared electromagnetic response of polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ [33]. The incoming mid-infrared light couples to the axion and the coupling creates far-infrared electromagnetic radiation at the resonance frequency $\omega_{\theta}$ of the transverse optical phonon that dominates the far-infrared response of the polar material. The mid-infrared response hits the axion resonance at frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ which is detuned by the frequency of the transverse optical phonon $\omega_{\theta}$ from the frequency $\omega_{A}=m_{A}c^{2}/\hbar=2\pi\times(122.7\pm 1.0)~{}{\rm THz}=2\pi c\times(4097\pm 32)~{}{\rm cm^{-1}}$ set by the axion rest-mass.
Measuring the mid-infrared frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ that resonantly excites the polar material provides a table-top determination of the mass of the axion. The size of the dip in mid-infrared transmission through a sample of the polar material on resonance at $\omega_{A\theta}$ is set by the infrared absorption coefficient [34]:
@@
where $\alpha(\omega_{\theta})$ is the absorption coefficient of the phonon, $\Gamma_{A\theta}\approx 2\pi c/t$ gives the width of the axion resonance in terms of the thickness of the sample $t$, and $\Gamma_{\theta}$ is the width of the phonon resonance. Using typical parameters for a polar material such as silicon nitride ${\rm Si_{3}N_{4}}$, we find the mid-infrared absorption coefficient at $\omega_{A\theta}$ from the axion resonance is large enough to be readily observable with samples that show a dip in the far-infrared at frequency $\omega_{\theta}$ from the transverse optical phonon [33]
@@
The mid-infrared absorption in polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ at frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ due to axions from the local dark matter halo of the galaxy can thus be seen with readily available samples and infrared spectrometers [33]
@@
Acknowledgements— This work was performed in part at the Kavli Institute for Theoretical Physics, which is supported by the National Science Foundation grant No. PHY-1748958, at the Aspen Center for Physics, which is supported by National Science Foundation grant No. PHY-2210452, at the Institute for Nuclear Theory, which is supported by the Department of Energy under grant No. DE-FG02- 00ER41132, and at the Mainz Institute of Theoretical Physics, which is part of the Cluster of Excellence PRISMA+ Project ID 39083149.

### 15 Mar 2024 v37

@@ Abstract
Recent measurements of muon spin precession confirm a long-standing tension with the standard model of particle physics. We argue that axions form the dark matter in the universe and that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument gives a percent level prediction for the mass of the axion that can be tested by measuring the mid-infrared electrodynamic response of a wide range of polar materials. A broad resonance in the scattered visible light intensity from quartz crystal at room-temperature and ambient pressure is reported that peaks at a frequency shift of the light which is consistent with the creation of axions with the predicted mass.
@@
$A_{M}$ $=$ $M_{L}^{-}\overline{M}_{R}^{-}-M_{R}^{+}\overline{M}_{L}^{+}-M_{L}^{+}\overline{M}_{R}^{+}+M_{R}^{-}\overline{M}_{L}^{-}$ (1)
@@
Next, we suppose that axions $A_{M}$ of each kind and photons $\gamma_{H}$ of each helicity formed in equal numbers in the early universe prior to baryogenesis. This fixes the ratio $n_{A}/n_{\gamma}=|M|/|H|=6$ of axions to photons and gives the axion rest-mass energy [8]:
$m_{A}c^{2}=2.70~{}kT_{\gamma}\left(\frac{n_{A}}{n_{\gamma}}\right)^{-1}\frac{\Omega_{A}h^{2}}{\Omega_{\gamma}h^{2}}=(0.508\pm 0.004)~{}{\rm eV},$ (2)
where, $T_{\gamma}=(2.7255\pm 0.0006)~{}{\rm K}$ is the present photon temperature [9], $\Omega_{\gamma}h^{2}=(2.473\pm 0.002)\times 10^{-5}$ is the photon energy density parameter and $\Omega_{A}h^{2}=0.11882\pm 0.00086$ is that of dark matter [10].
@@
Muon Haloscopy—Figure 1 shows how axions in the local dark matter halo shift muon spin precession [3, 4, 5]. Acting as a background field $\tilde{\phi}_{A}(q_{A})$ with on-shell four-momentum $q_{A}^{2}=(m_{A}c)^{2}\ll(m_{\mu}c)^{2}$, the axion $A$ couples to the virtual photon dressing the vertex of the positive muon $\mu^{+}$ and the magnetic field $B$. The axion-photon coupling of strength $g_{A\gamma\gamma}=(0.68\pm 0.02)\times 10^{-10}~{}{\rm GeV^{-1}}$ [12] changes the photon’s Green’s function [13, 14]
$\Delta D_{+}(x,x^{\prime})=m_{A}^{2}c^{4}g_{A\gamma\gamma}\tilde{\phi}(q_{A})D_{+}(x,x^{\prime}),$ (3)
@@
$\frac{\Delta a_{\mu}}{a_{\mu}}$ $=$ $m_{A}^{2}c^{4}g_{A\gamma\gamma}\tilde{\phi}_{A}(q_{A})$ (4)
$=$ ±$\pm m_{A}c^{2}g_{A\gamma\gamma}\sqrt{\rho_{A}N_{\mu}V_{\mu}T_{\mu}/\hbar}$
where, the axion field strength $|\tilde{\phi}_{A}(q_{A})|^{2}\approx\phi_{A}^{2}N_{\mu}V_{\mu}T_{\mu}/\hbar$ is fixed by the axion energy density $\rho_{A}=\phi_{A}^{2}m_{A}^{2}c^{4}$, the volume $V_{\mu}$ of the muon beam, the lifetime $T_{\mu}$ of the muon in the axion rest-frame, and the average number of muons $N_{\mu}$ during spin precession. The best astrophysical estimate for the energy density of dark matter in the local halo is $\rho_{A}=(0.30\pm 0.03)~{}{\rm GeV/cm^{3}}$ [16].
@@
$\Delta a_{\mu}$ $=$ $(171\pm 10)\times 10^{-11}\left(\frac{N_{\mu}}{1930}\right)^{1/2}\left(\frac{\sigma_{x}}{1.78~{}{\rm cm}}\right)^{1/2}$
@@
where, $\sigma_{x}$ is the width of the muon beam, $\sigma_{y}$ is its height, and $R$ is the radius of the storage ring. Combining the shifts found from Eq. (Weighing the Axion with Muon Haloscopy) using the muon beam parameters for the two experiments (FNAL E989 [2, 3] and BNL E821[4]) that dominate the experimental world average $a_{\mu}({\rm Exp})$, we find the weighted average shift
$\Delta a_{\mu}({\rm Exp})=(171\pm 10)\times 10^{-11},$ (6)
where, we have used the same weights for $\Delta a_{\mu}({\rm Exp})$ as the weights entering $a_{\mu}({\rm Exp})$ [18]. This shift roughly resolves the $3.4~{}\sigma$ tension between $a_{\mu}({\rm Exp})$ and the standard model value $a_{\mu}({\rm SM})$[19, 20, 21, 22, 23, 24]:
@@
where, we have used the recently emerged consensus value $a_{\mu}^{\rm HLO}({\rm SM})=(704.1\pm 3.0)\times 10^{-10}$ [25] for the leading order hadronic contribution to the muon spin precession frequency within the Standard Model [26].
A remarkable consequence of this resolution of the tension is that it can be tested in a relatively straightforward fashion. Flipping the magnetic storage ring over the muon beamline, such that the muons circulate with the opposite sense around the ring, would yield a spin precession frequency in tension with the present experimental world average by $2\Delta a_{\mu}({\rm Exp})/a_{\mu}({\rm Exp})=2\times(2.71\pm 0.17)~{}{\rm BNL}=(5.4\pm 0.3)~{}{\rm BNL}$, where $1~{}{\rm BNL}=0.54~{}{\rm ppm}$ [26] is the precision achieved at BNL E821 [4] and is the precision that can be achieved at FNAL E989 [2, 3] in a data collecting run with the flipped ring lasting roughly one month. The uncertainty of the present world average, roughly $0.4~{}{\rm BNL}$ [2, 26], combines with the statistics-dominated uncertainty of such a short run to give an expected significance of the tension $(5.1\pm 0.3)\sigma$ that surpasses the conventional threshold for the discovery of a new phenomenon in particle physics, where $\sigma\approx 1~{}{\rm BNL}$ is the uncertainty of the comparison.
Making Axions in Polar Materials with Infrared— Axions from the local dark matter halo of the galaxy can be seen in the mid-infrared electromagnetic response of a wide range of polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ [27]. The incoming mid-infrared light with frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ makes axions within the polar material with energy $\hbar\omega_{A}=m_{A}c^{2}=h\times(122.7\pm 1.0)~{}{\rm THz}$ by absorbing a transverse optical phonon at the far-infrared frequency $\omega_{\theta}\approx 2.4~{}{\rm THz}$ [27]. The axions leave the material and create a dip in the infrared transmission through the sample at the axion resonance frequency $\omega_{A\theta}$.
Measuring the mid-infrared frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ at which the polar material makes axions and the far-infared frequency $\omega_{\theta}$ of the transverse optical phonons, one finds the rest-mass energy of the axion $m_{A}c^{2}=\hbar\omega_{A}=\hbar(\omega_{A\theta}+\omega_{\theta}).$ The size of the dip in mid-infrared transmission through a sample of the polar material on the axion resonance at $\omega_{A\theta}$ is set by the infrared absorption coefficient [28]:
@@
where $\alpha(\omega_{\theta})$ is the absorption coefficient of the phonon, $\Gamma_{A\theta}\approx 2\pi c/t$ gives the width of the axion resonance in terms of the thickness of the sample $t$, and $\Gamma_{\theta}$ is the width of the phonon resonance.
Using typical parameters for a polar material such as silicon nitride ${\rm Si_{3}N_{4}}$, we find the mid-infrared absorption coefficient at $\omega_{A\theta}$ from the axion resonance [27]
@@
The mid-infrared absorption with coefficient $\alpha(\omega_{A\theta})\approx 4\alpha(\omega_{\theta})$ in polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ at frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ due to creation of axions can thus be seen with readily available samples and infrared spectrometers provided the spectrometer sees the dip with absorption coefficient $\alpha(\omega_{\theta})$ in the far-infrared transmission spectrum at frequency $\omega_{\theta}$ from the transverse optical phonon in the sample [27].
Figure 2: Tentative observation of the axion resonance in the Raman light scattering spectrum of synthetic quartz crystal (${\rm SiO_{2}}$, RRUFF deposition R110108) [30, 29] at room temperature and ambient pressure. The axion rest-mass (See Eq. 2) gives the window (vertical lines) for the frequency shift $\omega_{A}=m_{A}c^{2}/\hbar=2\pi c\times(4096\pm 32)~{}{\rm cm^{-1}}$ of visible light that makes axions within the polar material. A broad band of scattered light intensity is observed with the peak frequency shift falling within the predicted axion resonance window.
Axion Resonance in Quartz Raman Spectroscopy— A tentative observation of the axion resonance in the mid-infrared electrodynamics response of quartz (${\rm SiO_{2}}$) [29] has been reported by the RRUFF project [30] using Raman light scattering spectroscopy (Fig. 2). The scattered light intensity has a broad peak centered on the frequency shift $\omega_{A}({\rm SiO_{2}})=2\pi c\times(4080\pm 4)~{}{\rm cm^{-1}}$ which falls within the predicted window $\omega_{A}=m_{A}c^{2}/\hbar=2\pi c\times(4096\pm 32)~{}{\rm cm^{-1}}$. Quartz has no known Raman-active vibration modes in this frequency range and chemical analysis suggests that the sample has insufficient concentration of luminescent impurities to account for the strength of the observed resonance [29].
Assuming that visible light incident on the quartz crystal makes axions within the sample, the frequency shift of the scattered light where the intensity peaks $\omega_{A}({\rm SiO_{2}})$ can be interpreted in terms of the rest-mass energy of the axion:
$m_{A}({\rm SiO_{2}})c^{2}=\hbar\omega_{A}({\rm SiO_{2}})=(0.5059\pm 0.0005)~{}{\rm eV}.$ (10)
This agrees with the estimate in Eq. (2) but is roughly ten times more precise. If this interpretation is confirmed, then it would provide strong support for the novel physical picture of dark matter in terms of axions with number density $n_{A}=6n_{\gamma}$.
Indeed, the rest-mass energy of the axion from quartz in Eq. (10), when combined with the axion number density $n_{A}=6n_{\gamma}=(2464.2\pm 1.6)~{}{\rm cm^{-3}}$ [8] suggested by the novel physical picture of dark matter, gives the cosmological energy density parameter for axions:
$\Omega_{A}h^{2}({\rm SiO_{2}})=\frac{8\pi G}{3}\frac{h^{2}}{H_{0}^{2}}n_{A}m_{A}=(0.11831\pm 0.00008),$ (11)
where, $H_{0}/h=100~{}{\rm km/(sec\cdot Mpc)}=1/(9.778~{}{\rm Gyr})$ [8] gives the conventional parameterization of the present expansion rate of the universe. This agrees with the best estimate for the cosmological energy density parameter of dark matter $\Omega_{A}h^{2}=(0.11882\pm 0.00086)$ [10]. However, it is ten times times more precise.
@@
Acknowledgements— The author acknowledges the support of Daria Mazura.